OpenStudy (anonymous):

How do i solve for this question using synthetic division You are given f(x) = x^4 - 3x^3 - kx - 5. When f(x) is divided by x + 3, the remainder is 166

jimthompson5910 (jim_thompson5910):

By the remainder theorem, you can say that f(-3) = 166

jimthompson5910 (jim_thompson5910):

f(x) = x^4 - 3x^3 - kx - 5 f(-3) = (-3)^4 - 3(-3)^3 - k(-3) - 5 ... plug in x = -3 166 = (-3)^4 - 3(-3)^3 - k(-3) - 5 ... plug in f(-3) = 166 now solve for k

OpenStudy (anonymous):

thank you so much

jimthompson5910 (jim_thompson5910):

you're welcome